OpenStudy (anonymous):

Teesha is in the French club. There are 31 students in the club. The French teacher will pick two students at random to guide visiting students from France. What is the probability that Teesha will be picked as a guide? Hint: Teesha must be picked first or second. 1/30 30/31 1/930 2/31

OpenStudy (anonymous):

picked first - P = 1/31 picked second - P = 30/31 * 1/30 = 30/930 = 1 / 31 required P = 1/31 + 1/31 = 2/31

OpenStudy (anonymous):

second line : - the probabilities are multiplied because they are independent last line they are added as they are dependent - its one OR the other